OpenStudy (anonymous):

Hi all. I have 4 problems that I am having a very hard time with so need some help please. Solve the system of equations by graphing: 7x-5y=15 & 7x=5y+15. Need solution set (ordered pair)

OpenStudy (anonymous):

do you know how to plot graphs?

OpenStudy (anonymous):

I need feedback to help you.

OpenStudy (anonymous):

sorry had to step away. I have a hard time with graphing.

OpenStudy (anonymous):

okay. graphs have a dependent variable and an independent variable. The x axis represents the independent variable and the y axis represents the independent variable.

OpenStudy (anonymous):

so, now choose any values of x for your equation 7x - 5y = 15. That is, choose any values of x and you will get a corresponding dependent value of y.

OpenStudy (anonymous):

you said the x and y both represent the independent.

OpenStudy (anonymous):

sorry, y represents the dependent variable.

OpenStudy (anonymous):

let me give you an example. suppose you are a salesman. Your year end bonus is 10% of sales you make that year. So, the sales you make is an independent variable. You can make 10000 dollars worth of sales, or if you are better than average, you make 20000 dollars of sales. Or if you are brilliant, you make $100,000 of sales. But notice what happens. The bonus is 10% of your sales. So the more sales you make, the more bonus you get. In other words, the bonus is dependent on sales.

OpenStudy (anonymous):

unfortunately all dependent variables are not so easily represented as 10% of the independent variables. suppose your bonus was 10% of sales - medicare - 1/2taxes

OpenStudy (anonymous):

then your equation would be bonus = sales/10-M-1/2tax

OpenStudy (anonymous):

anyway, the point is, there are independent variables which are in your hand.

OpenStudy (anonymous):

and there are other variables which depend on your independent variable.

OpenStudy (anonymous):

so in your equation, 7x - 5y = 15, let say x is the independent variable, just because it is nice to have the x variable on the independent axis (x-axis)

OpenStudy (anonymous):

so then, what is the dependent variable, in terms of x?

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

5y?

OpenStudy (anonymous):

if bonus is 10% of sales, express bonus in terms of sales. similarly if 7x-5y is 15, express y in terms of x.

OpenStudy (anonymous):

basically, you are trying to express the dependent variable as a function of the independent variable.

OpenStudy (anonymous):

you could say that 7x-5y = 15 is " 7 times your sales minus 5 times your bonus always gives 15 grand"

OpenStudy (anonymous):

so express bonus in english, or in an equation.

OpenStudy (anonymous):

are you getting what I am saying?

OpenStudy (anonymous):

ah not really

OpenStudy (anonymous):

sorry that is one area I just seem can' get a grasp on

OpenStudy (anonymous):

okay if 7 times sales minus 5 times bonus is always 15 grand, then 7 times sales minus 15 grand divided by 5 is the bonus. in other words, y = (7x-15)/5

OpenStudy (anonymous):

so choose your x values and find out the corresponding y values.

OpenStudy (anonymous):

since this is the equation for a line, you need only two points.

OpenStudy (anonymous):

ok so here is what I did. tell me if I am rght

OpenStudy (anonymous):

I chose 2 for x. y=7(2)-15/5.... y=14-15/5.... y=-1/5

OpenStudy (anonymous):

so how did I do?

OpenStudy (anonymous):

okay thats great!

OpenStudy (anonymous):

but 1/5 is kinda hard to plot on a graph.

OpenStudy (anonymous):

yes very true. do I need to choose a different number or x?

OpenStudy (anonymous):

notice that y i some value/5. so if the value is a multiple of 5, then you would get a nice round number for y.

OpenStudy (anonymous):

y is value/5.

OpenStudy (anonymous):

so suppose you chose x = 5. 7 times 5 is 35, and 35-15 is 20. 20 is a multiple of 5. so y =4.

OpenStudy (anonymous):

so one value could be (5,4)

OpenStudy (anonymous):

ok I changed the value of x to 5. I came up with y= -3

OpenStudy (anonymous):

both are nice round numbers, easy to plot on a pice of graph paper.

OpenStudy (anonymous):

how did you come up with y = -3 for x = 5?

OpenStudy (anonymous):

sorry fiured wrong. redid it and got 4.

OpenStudy (anonymous):

okay, next choose x =10

OpenStudy (anonymous):

ok I got 11

OpenStudy (anonymous):

okay great! so you have two points (5,4) and (10,11). there is only one line which passes through both points. So you have your first line down on the graph! huzzah! one down, one equation to go.

OpenStudy (anonymous):

right

OpenStudy (anonymous):

so what is the second equation?

OpenStudy (anonymous):

7x=5y+15

OpenStudy (anonymous):

this is the same equation as before.

OpenStudy (anonymous):

the other one was 7x-5y=15

OpenStudy (anonymous):

okay, subtract 5y from both sides of your second equation and post what you got.

OpenStudy (anonymous):

y=-7x+15

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

how did you get that? http://www.dabbleboard.com/draw?b=Guest644716&i=0&c=3327a9e1158d904a38fbad735c24ed28cad7bcd9 show me there.